The heights of 1,000 college students are normally distributed about a mean value of 170 cm. The standard deviation is 7.2 cm. What is the probability of selecting a students smaller than 158 centimeters.


1
Expert's answer
2022-03-24T19:55:28-0400

"\mu=170,\sigma=7.2"

"P(X<158)=P(Z<\frac{158-170}{7.2})"

"=P(Z<-1.67)"

From the normal tables

"=0.0475"